OpenStudy (anonymous):

I Need Help With 3p^2 - 2p - 5

OpenStudy (anonymous):

3p^2 - 2p - 5 3p^2 + 3p -5p -5 3p(p + 1) -5(p + 1) (3p - 5) (p + 1) 3p - 5 = 0 3p = 5 p = 5/3 p + 1 =0 p = -1 So, p=5/3 and p=-1 are the two solution.

OpenStudy (anonymous):

what do u want to find?